OpenStudy (neverendingcycle):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 A. 3x – 8y = 5 B. 8x + 3y = 5 C. 8x – 3y = 5 D. 8x – 3y = –5

OpenStudy (neverendingcycle):

Thanks! Are you sure?

OpenStudy (anonymous):

www.geteasysolution.com/7-3(x-y)=5x+2

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

I'm 100% sure it's B

OpenStudy (neverendingcycle):

Not d?

OpenStudy (anonymous):

No it's B

OpenStudy (anonymous):

Look at the website :)

OpenStudy (neverendingcycle):

okay then....If i get this wrong you suck! JK XD

OpenStudy (anonymous):

LOL Ok

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

What if I already suck jk

OpenStudy (neverendingcycle):

XD

OpenStudy (anonymous):

hold up... we have - 8x + 3y = -5 now first term must be positive so multiply -1

OpenStudy (anonymous):

www.geteasysolution.com/7-3(x-y)=5x+2

OpenStudy (anonymous):

check this website

OpenStudy (neverendingcycle):

WBRAAARRRG

OpenStudy (anonymous):

What

OpenStudy (neverendingcycle):

WHOS RIGHT THERE ARE TO MANY DIFFERENT ANSWERS!

OpenStudy (anonymous):

It's B trust me

OpenStudy (neverendingcycle):

okay then

OpenStudy (anonymous):

I know it is

OpenStudy (anonymous):

no B is wrong

OpenStudy (anonymous):

nope

OpenStudy (neverendingcycle):

STOP CONFUSING ME!! WHAT IS THE ANSWER?

OpenStudy (anonymous):

its either C or D.... B to me seems way off

OpenStudy (neverendingcycle):

GHFCCFGVBHJRVHBKCDJCFKVGKB

OpenStudy (anonymous):

7-3 (x-y) = 5x+2 *you distribute -3 to (x-y) 7-3x+3y = 5x+2 *then group like terms -3x-5x+3y=2-7 -8x+3y=-5 *then you multiply everything by -1 since u want to have a positive x (-8x+3y=-5)-1 8x-3y=5

OpenStudy (neverendingcycle):

WHICH ONE?!!!!! XD

OpenStudy (anonymous):

@Directrix am i correct?

OpenStudy (neverendingcycle):

I hope

OpenStudy (anonymous):

7 + -3(x + -1y) = 5x + 2 7 + (x * -3 + -1y * -3) = 5x + 2 7 + (-3x + 3y) = 5x + 2 Reorder the terms: 7 + -3x + 3y = 2 + 5x Solving 7 + -3x + 3y = 2 + 5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. 7 + -3x + -5x + 3y = 2 + 5x + -5x Combine like terms: -3x + -5x = -8x 7 + -8x + 3y = 2 + 5x + -5x Combine like terms: 5x + -5x = 0 7 + -8x + 3y = 2 + 0 7 + -8x + 3y = 2 Add '-7' to each side of the equation. 7 + -8x + -7 + 3y = 2 + -7 Reorder the terms: 7 + -7 + -8x + 3y = 2 + -7 Combine like terms: 7 + -7 = 0 0 + -8x + 3y = 2 + -7 -8x + 3y = 2 + -7 Combine like terms: 2 + -7 = -5 -8x + 3y = -5

OpenStudy (anonymous):

7 - 3 ( x - y ) = 5x + 2 -7 -7 --------------------- -3x + 3y = 5x -5 -5x -5x --------------- -8x + 3y = -5 next step must multiply by (-1) to get answer

OpenStudy (anonymous):

Add '-3y' to each side of the equation. -8x + 3y + -3y = -5 + -3y Combine like terms: 3y + -3y = 0 -8x + 0 = -5 + -3y -8x = -5 + -3y

OpenStudy (anonymous):

@NeverEndingCycle when do u know the answer

OpenStudy (anonymous):

Ax + By = C ..............Ax term must be positive to be in standard form

OpenStudy (anonymous):

when you multiply by (-1) (-8x +3y ) = (-1) -5 you get the answer

OpenStudy (anonymous):

but the + stays the same

OpenStudy (anonymous):

standard form can have - By or - C just not the first term

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

wahtever

OpenStudy (neverendingcycle):

im so cunfuzled

OpenStudy (neverendingcycle):

what is the answer?

OpenStudy (anonymous):

you had it correct to the point where -Ax and to be in standard form you need to multiply both sides since the equation is already in position of standard form

OpenStudy (neverendingcycle):

so the answer is.......................?

OpenStudy (neverendingcycle):

a b c or d

OpenStudy (neverendingcycle):

?

OpenStudy (anonymous):

answer is